OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 28 vowels and 16 consonats, what is the probability you will choose a vowel then a consonat ?

OpenStudy (mathstudent55):

There are 28 vowels and 16 consonants. That means there are 28 + 16 = 44 tiles Since the first tile is replaced these are two independent events. P(vowel then consonant) = P(vowel) * P(consonant) P(vowel) = 28/44 P(consonant) = 16/44